OpenStudy (anonymous):

Help with? 7-3x^2=(x+2)^2-6 I can't find the solutions..

OpenStudy (anonymous):

the equation -4x^2-4x+9=0; so use the quadratic formula...and u'll get the answers as, x=(4+4*sqrt(10))/-2*4 and x=x=(4-4*sqrt(10))/-2*4 u know what and how to use quadratic formula ?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

so 3/5 and 2/3

OpenStudy (anonymous):

ok first post ur options...please

OpenStudy (anonymous):

1/5, 1/3, 3/5, 1/2, 2/3, and 4/5.

OpenStudy (anonymous):

7-3x^2 = x^2+4x+4-6 Get all terms on one side 4x^2+4x-9=0 Solve using graphing calculator, and finding where y = 0 x=1.081 or x= -2.081

OpenStudy (anonymous):

the final simplified equation is x^2+x-9/4=0 u'll find that none of the choices satisfy...believe me . the answer i gav earlier is right...u can cross check with anyone..

OpenStudy (anonymous):

7-3x^2=x^2 +4x +4 -6 7-3x^2 = x^2 +4x -2 0 =4x^2+4x-9 now use the formula x=-b+ or - sqr root(b^2 - 4 ac)/2a a=4, b=+4, c=-9 x=--4+-sqr root(16-4(4)(-9)/2(4) =+4+-sqr root(16+144)/8 =(4+-sqr root (160))/8 =4/8 +- (sqr root 160)/8 =1/2 +- 1/2(sqr root 10)

OpenStudy (anonymous):

so shaunlampert also got the same answer :)

OpenStudy (anonymous):

because of the sqr root 10 the number cannot be represented as a fraction. Hence none of the choices satifies the equation as suggested by brackett